OpenStudy (anonymous):

find integral of (xcube-1)raiseto1/3*xraiseto5

OpenStudy (anonymous):

( x^3 - 1) / 3 x^5

OpenStudy (anonymous):

is this d Q?

OpenStudy (anonymous):

how did u do it?!?

OpenStudy (anonymous):

i didnt solve yet...i have written d integrand again...is it right?

OpenStudy (anonymous):

exactly i dont know ...thats why im asking u how u did it

OpenStudy (anonymous):

(xcube-1)raiseto1/3*xraiseto5 = ( x^3 - 1) / 3 x^5 ????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thts the question

OpenStudy (anonymous):

( x^3 - 1) / 3 x^5 = 1/3 x^(-2) - 3 x^(-5)

OpenStudy (anonymous):

1/3 x^(-2) - 1/3 x^(-5)

OpenStudy (anonymous):

theres one more.. fine integral of 1+cosAcosx/cosA+cosx ....(A is a constant!)

OpenStudy (anonymous):

(1+cosAcosx)/ (cosA+cosx) ???

OpenStudy (anonymous):

1/3 is raise to that whole x^3-1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(1/3) x^(-2) - (1/3 )x^(-5) but it's not d answer...u have to integrate it using integration of x^n = x^(n+1)/(n+1)

OpenStudy (anonymous):

k..

OpenStudy (anonymous):

is it substitution,by parts or what?!

OpenStudy (lgbasallote):

\[\Large \int (x^3 - 1)^{\frac 13} x^5 dx\] let \[u = x^3 - 1\] \[du = 3x^2dx\] so the integral becomes \[\Large \implies \frac 13 \int u^{\frac 13} du \times (u-1)\] so if you simplify... \[\Large \implies \frac 13 \int (u^{\frac 43} - u ^{\frac 13})du\]

OpenStudy (lgbasallote):

uhh wait..that should be + not -

OpenStudy (anonymous):

(xcube-1)raiseto1/3*xraiseto5

OpenStudy (anonymous):

wait...check d question again..

OpenStudy (anonymous):

there is no 1/3 rd power ...

OpenStudy (anonymous):

ill check

OpenStudy (lgbasallote):

raise to 1/3 means ^1/3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh...yes...sorry

OpenStudy (anonymous):

question is right

OpenStudy (lgbasallote):

anyway...like i was saying..should be + not - in my last line \[\huge \implies \frac 13 \int (u^{\frac 43} + u^{\frac 13})du\]

OpenStudy (anonymous):

yes...lgbasallote is right

OpenStudy (lgbasallote):

now integrate that

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thx

OpenStudy (lgbasallote):

welcome